o) Calculate the loonds cureakipyled? b) Colculale the bands VTM and Cucrant viei in parst ey Used 뎌 Ganeci son-afile bonds, current Poe. and alu ex plein
Answer #1

a) Current Yield = Coupon/Current Price = 11%*1000/1150 = 10.48%

b) YTM using financial calculator

PMT = 1000*11% = 100

N = 18; PMT = 100; FV = 1000; PV = -1150; CPT I/Y

YTM = 9.26%

c) YTM is 9.26%, Current Yield is 10.48%, Coupon Rate = 11%

Since the Price of bond is more than par value hence YTM<Current Yield< Coupon Rate.

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well